# Stop-and-go waves reconstruction via iterative refinement

### v1

@@ Abstract
Stop-and-go waves are a fundamental phenomenon in freeway traffic flow, contributing to inefficiencies, crashes, and emissions. Recent advancements in high-fidelity sensor technologies have improved the ability to capture detailed traffic dynamics, yet such systems remain scarce and costly. In contrast, conventional traffic sensors are widely deployed but suffer from relatively coarse-grain data resolution, potentially impeding accurate analysis of stop-and-go waves. This article explores whether generative AI models can enhance the resolution of conventional traffic sensor to approximate the quality of high-fidelity observations. We present a novel approach using a conditional diffusion denoising model, designed to reconstruct fine-grained traffic speed field from radar-based conventional sensors via iterative refinement. We introduce a new dataset, I24-WaveX, comprising 132 hours of data from both low and high-fidelity sensor systems, totaling over 2 million vehicle miles traveled. Our approach leverages this dataset to formulate the traffic measurement enhancement problem as a spatio-temporal super-resolution task. We demonstrate that our model can effectively reproduce the patterns of stop-and-go waves, achieving high accuracy in capturing these critical traffic dynamics. Our results show promising advancements in traffic data enhancement, offering a cost-effective way to leverage existing low spatio-temporal resolution sensor networks for improved traffic analysis and management. We also open-sourced our trained model and code to facilitate further research and applications.
@@ 1 Introduction
Stop-and-go waves are among the most ubiquitous and significant phenomena in freeway traffic flow research[1; 2; 3]. Stop-and-go waves have been recognized as a major contributor to the system inefficiencies, crashes [4] and emissions [5; 6].
Modern freeway traffic management centers (TMCs) commonly utilize sensors such as loop detectors [7] and radar detectors [8] in daily operations to implement various corridor management strategies [9]. These sensors typically provide aggregated traffic measurements at typical intervals of 30 seconds to 5 minutes and are usually deployed at 0.5 to 1-mile intervals along the freeway. Because of such sensors, we know that stop-and-go traffic waves are extremely prevalent in freeway corridors [10]. However, while radar and inductive loop sensors offer a good cost-benefit ratio compared to expensive high-fidelity sensors for TMCs, their lack of fine spatio-temporal resolution means that they cannot accurately measure periods of low-speed travel temporally close to periods of high-speed travel [11; 12]. This underestimation is particularly problematic as these low-speed parts are critical for accurately estimating stop-and-go wave characteristics.
In recent years, advanced sensor technologies such as drones [13; 14], multi-camera monitoring systems [15; 16], LiDAR [17; 18], and on-board vehicle perception systems [19] have allowed the collection of precise vehicle trajectories, which has enabled rapid advancement in the analysis and modeling of stop-and-go waves [20; 21; 22; 16]. High-fidelity traffic measurements have demonstrated the capability to capture the intricate patterns of stop-and-go waves, which exhibit considerable complexity, as shown in [23; 24]. Figure 1 demonstrates a comparison of the traffic speed profiles observed from a conventional sensor system and high-fidelity sensor systems. Unfortunately, very few roadways worldwide currently have access to high-fidelity traffic instruments mentioned above capable of continuously collecting extensive vehicle trajectory data [25]. This significantly limits the generalizability and scalability of the observation and analysis of stop-and-go traffic to a broader context.
@@
Recent years have seen a flurry of advancements in generative artificial intelligence (AI) models, and their diverse applications [26; 27] have leveraged the potential of large-scale data via generative models, motivating the primary thesis of this work: generative models can learn the underlying mechanisms of traffic from low spatio-temporal resolution sensor input to generate enhanced data which better capturing stop-and-go waves, almost to the resolution of high-fidelity traffic sensors.
To investigate this, we collected 132 hours of rush-hour traffic data over 49 days from a stretch of 4 miles of roadway, amounting to over 2 million vehicle miles traveled (VMT). The dataset, named I24-WaveX, was gathered simultaneously from two distinct sensor systems: one from a commonly used freeway radar detector system (with sensors deployed every 0.3 to 0.5 mile and aggregated every 30 seconds), and the other from a camera-based traffic testbed capable of capturing individual vehicle trajectories (aggregated into speed profile data with bin resolution of 4 seconds and 0.02 miles, approximately 32 meters) [16; 28]. We then formulated the data enhancement problem as a traffic measurement super-resolution reconstruction task, with the goal of reconstructing the higher-fidelity data using only the lower fidelity data as input. To this end, we implemented a conditional diffusion denoising model for iterative refinement. Finally, we evaluated the extent to which our proposed method accurately reconstructs the details of stop-and-go waves.
@@
  1. (1)
    We are the first to apply a generative conditional diffusion denoising model designed to improve the spatio-temporal resolution of traditional freeway roadside traffic sensors, addressing the traffic measurement super-resolution challenge.
  2. (2)
    We produce the I24-WaveX dataset, a new large-scale empirical dataset over 2 million vehicle miles traveled featuring stop-and-go traffic, which was collected concurrently using both high-resolution and low-resolution sensor systems over 49 days, 4 to 4.5 hours a day. This dataset is intended to investigate the potential of generative models in analyzing traffic flow.
  3. (3)
@@
  4. (4)
    We release the code and fine-trained model parameters trained on the data to enhance the reproduciblilty of this paper future research.
The rest of this article is organized as follows: first, we briefly discuss related work on traffic waves and applications of generative models to traffic. We then describe the method proposed in this work, Next we illustrate the data collected and the experiments used to validate our methods, and present the obtained results. The article concludes with a discussion on the ability of the method to capture fine-grained wave properties and its limitations.
@@ 2 Related work and challenges
Traffic reconstruction across all scales [29] requires data collection as the first step. Efforts in empirical traffic data collection have been ongoing since 1933 [30], and these efforts can be categorized into microscopic and macroscopic scales. Macroscopic traffic data [30; 31; 32; 10; 33; 8], which refers to the aggregated traffic measurements typically collected from loop detectors, provides insights into collective phenomena such as the fundamental diagram and capacity drop [34; 35] On the other hand, microscopic traffic data such as vehicle trajectory data in a common reference system[36; 21], helps researchers understand individual vehicle behaviors.
@@
Traffic data which captures observations of both microscopic and macroscopic phenomena has significant value in the transportation research community. A relative scarcity of these data has limited empirical validation of theory in individual dynamics and traffic system dynamics. To date, only a few datasets are available which capture both microscopic and macroscopic data. Data collected from mobile and fixed sensors by the Mobile Century experiments [37] have been utilized for traffic reconstruction. Additionally, INRIX and loop detector data [12] have been collected and employed for cross-validation purposes. Purveyors of GPS-navigation applications have vehicle trajectory and aggregate data, but these are proprietary. Further efforts and contributions are needed to expand the availability and diversity of datasets in this field.
@@
Ideally, microscopic and macroscopic scales in traffic measurement could be converted into another, similar to the concept of downsampling and upsampling in image processing [38]. In practice, transitioning from microscopic to macroscopic is straightforward through aggregation via Edie’s definition [2], whereas transforming from macroscopic to microscopic is more challenging due to the complex and heterogeneous dynamics of individual particles in the traffic flow system. The problem is often referred to as traffic reconstruction [37; 39; 40]. One paradigm from previous research involves calibrating sets of microscopic modelled driving behaviors to replicate macroscopic measurements or phenomena [41]. A limitation of using micro-simulation to reproduce macroscopic phenomena is the inherent mismatch between the model and real-world aggregated data, which may not fully capture the complex and dynamic nature of actual traffic systems [42; 39]. Another paradigm involves using kernel smoothing [43; 44] and similar techniques to interpolate and impute sparse observations, in order to enhance the resolution and completeness of traffic data by filling in gaps and smoothing out inconsistencies. The adaptive smoothing method [45; 43; 46] is a widely recognized approach in this direction. More recently, the refinement problem was proposed and defined by [47], which utilized a simple linear regression approach that accounts for different traffic states to enhance traffic measurement resolution. The method proposed in that study achieves a 2 by 2 lift in resolution for both space and time each time, demonstrating the potential to improve resolution from 200 meters by 60 seconds to 50 meters by 30 seconds. In other fluid dynamical systems, deep learning has been explored and applied to the reconstruction problem. These approaches have shown promise in accurately capturing and reconstructing complex fluid dynamics [48; 49; 50], suggesting potential applicability to traffic flow systems [51] as well.
@@
In recent years, the field of image processing has seen significant advancements due to generative artificial intelligence. Generative models have enabled the transformation of low-resolution images into high-resolution ones via a process known as super-resolution [52]. This aligns well with the problem of enhancing traffic measurement resolution. The concept of generative deep learning gained prominence with the introduction of generative adversarial networks (GANs) [53], which have been extensively applied in traffic systems for tasks such as traffic data imputation [54; 55] and traffic state estimation [56]. More recently, diffusion probabilistic models have emerged as a promising approach [27]. These models have been investigated for tasks such as noisy data recovery [57] and GPS trajectory data generation [58], demonstrating their robustness to hyperparameters and their potential for a wide range of applications. These advancements in generative diffusion models offer new opportunities for improving traffic measurement resolution and addressing the challenges associated with low spatio-temporal resolution data, which have not been investigated before.
@@ 3.1 Problem formulation
In this article, stop-and-go wave enhancement refers to the problem of recovering fine-grained speed data (high-fidelity) from corresponding coarse-grained traffic measurement data (relatively low-fidelity), which can be viewed as a super-resolution task. Coarse-grained data typically collected from radar detectors or loop detectors, is considered low spatio-temporal resolution data and is denoted by $\mathbf{r}\in\mathcal{R}$. Fine-grained speed data, generated from vehicle trajectory data, is viewed as high-fidelity data and is represented by $\mathbf{m}\in\mathcal{M}$. For a given space and time range, forming tensors $\mathbf{r}$ and $\mathbf{m}$ herein, each observation of coarse-grained data is represented as $\mathbf{r}\in\mathbb{R}^{T_{r}\times S_{r}\times 3}$ and fine-grained data is represented as $\mathbf{m}\in\mathbb{R}^{T_{m}\times S_{m}\times 1}$, where $T_{r},~{}S_{r},~{}T_{m},~{}S_{m}$ are the number of time and space observations in the space-time range, respectively. Typically, $T_{r}<T_{m}$ and $S_{r}<S_{m}$, which induces a dimension lifting in the output. The coarse-grained data contains three traffic measurements: speed, volume, and occupancy, whereas the fine-grained data includes only one measurement: speed. The problem addressed in this context is to find a mapping function $f:\mathcal{R}\to\mathcal{M}$ that can restore the fine-scale details of stop-and-go waves in traffic from the coarse-grained data.
@@ 3.2 Conditional denoising diffusion model
We adapt the denoising diffusion probabilistic model (DDPM) [59] for conditional fine-grained speed field generation. The method contains a noising process (e.g. forward diffusion) and a denoising process (e.g. inverse diffusion) as shown in Figure 2. These processes are defined as:
@@
$q(\mathbf{m}_{1:T}\mid\mathbf{m}_{0})$ $=\prod_{t=1}^{T}q(\mathbf{m}_{t}\mid\mathbf{m}_{t-1})$ (1)
$q(\mathbf{m}_{t}\mid\mathbf{m}_{t-1})$ $=\mathcal{N}(\mathbf{m}_{t}\mid\sqrt{\alpha_{t}}\mathbf{m}_{t-1},(1-\alpha_{t})\mathbf{I})$ (2)
@@
where $\gamma_{t}=\prod_{t=1}^{t}\alpha_{i}$.
Learn the inverse diffusion (denoise): to reverse the forward diffusion process, a neural denoising model $f_{\theta}$ takes the coarse-grained data $\mathbf{r}$ and the noisy target image $\mathbf{\tilde{m}}$ as defined in Equation 3,
$\mathbf{\tilde{m}}=\sqrt{\gamma}\mathbf{m}_{0}+\sqrt{1-\gamma}\epsilon,~{}~{}~{}\epsilon\in\mathcal{N}(0,\mathbf{I})$ (4)
The model $f_{\theta}(\mathbf{r},~{}\mathbf{\tilde{m}},~{}\gamma)$ is trying to learn from the coarse data $\mathbf{r}$, the noisy target image $\mathbf{\tilde{m}}$, as well as the variance of the noise $\gamma$, to predict the noise vector $\epsilon$ in Equation 4. Here, $\gamma$ is the scalar parameter drawn from the distribution $p(\gamma)$. The objective function for training $f_{\theta}$ is written as:
@@
During the training process, each iteration takes a gradient descent step:
$\nabla_{\theta}\left\|f_{\theta}(\mathbf{r},\sqrt{\gamma}\mathbf{m}_{0}+\sqrt{1-\gamma}\epsilon,\gamma)-\epsilon\right\|_{2}^{2}$ (6)
until convergence. The detailed proof and justification of the choice of the objective function for $f_{\theta}$ from the perspective of a variational lower bound and denoising score-matching can be found in [27; 59].
@@
$\mathbf{W}_{ij}(\mathbf{r},v_{c})=\begin{cases}\omega&\text{if }\mathbf{m}_{ij}<v_{c}\\ 1-\omega&\text{if }\mathbf{m}_{ij}\geq v_{c}\end{cases},$ (7)
@@
where $z$ is defined as Gaussian noise for all the steps other than $t=1$ (i.e., $z\sim\mathcal{N}(0,\mathbf{I})\text{ if }t>1,\text{ else }z=0$). The derivation of Equation 9 can be found in [59].
Neural denosing model choice: the UNet [61] architecture is a popular choice for diffusion denoising model regression [27; 59] because its encoder-decoder structure with skip connections effectively captures multi-scale features and preserves high-resolution details throughout the denoising process. Detailed task specific architectural details can be found in the EXPERIMENTS section.

### v2

# Stop-and-go wave super-resolution reconstruction via iterative refinement

@@ Abstract
Stop-and-go waves are a fundamental phenomenon in freeway traffic flow, contributing to inefficiencies, crashes, and emissions. Recent advancements in high-fidelity sensor technologies have improved the ability to capture detailed traffic dynamics, yet such systems remain scarce and costly. In contrast, conventional traffic sensors are widely deployed but suffer from relatively coarse-grain data resolution, potentially impeding accurate analysis of stop-and-go waves. This article explores whether generative AI models can enhance the resolution of conventional traffic sensor to approximate the quality of high-fidelity observations. We present a novel approach using a conditional diffusion denoising model, designed to reconstruct fine-grained traffic speed field from radar-based conventional sensors via iterative refinement. We introduce a new dataset, WaveX, comprising 132 hours of data from both low and high-fidelity sensor systems, totaling over 2 million vehicle miles traveled. Our approach leverages this dataset to formulate the traffic measurement refinement problem as a spatio-temporal super-resolution task. We demonstrate that our model can effectively reproduce the patterns of stop-and-go waves, achieving high accuracy in capturing these critical traffic dynamics. Our results show promising advancements in traffic data refinement, offering a cost-effective way to leverage existing low spatio-temporal resolution sensor networks for improved traffic analysis and management. We also open-source our dataset, trained model and code to enable further research and applications.
@@ 1 Introduction
Stop-and-go waves are among the most ubiquitous and significant phenomena in freeway traffic flow research[1; 2; 3]. Stop-and-go waves have been recognized as a major contributor to the system inefficiencies, crashes [4] and emissions [5; 6]. Modern freeway traffic management centers (TMCs) commonly utilize sensors such as loop detectors [7] and radar detectors [8] in daily operations to implement various corridor management strategies [9]. These sensors typically provide aggregated traffic measurements at typical intervals of 30 seconds to 5 minutes and are usually deployed at 0.5 to 1-mile intervals along the freeway. Because of such sensors, we know that stop-and-go traffic waves are extremely prevalent in freeway corridors [10; 11]. However, while radar and inductive loop sensors offer a good cost-benefit ratio compared to expensive high-fidelity sensors for TMCs, their lack of fine spatio-temporal resolution means that they cannot accurately measure periods of low-speed travel temporally close to periods of high-speed travel [12; 13]. This underestimation is particularly problematic as these low-speed parts are critical for accurately estimating stop-and-go wave characteristics.
In recent years, advanced sensor technologies such as drones [14; 15], multi-camera monitoring systems [16; 17], LiDAR [18; 19], and on-board vehicle perception systems [20] have allowed the collection of precise vehicle trajectories, which has enabled rapid advancement in the analysis and modeling of stop-and-go waves [21; 22; 23; 17]. High-fidelity traffic measurements have demonstrated the capability to capture the intricate patterns of stop-and-go waves, which exhibit considerable complexity, as shown in [24; 25]. Figure 1 demonstrates a comparison of the traffic speed profiles observed from a conventional sensor system and high-fidelity sensor systems. Unfortunately, very few roadways worldwide currently have access to high-fidelity traffic instruments mentioned above capable of continuously collecting extensive vehicle trajectory data [26]. This significantly limits the generalizability and scalability of the observation and analysis of stop-and-go traffic to a broader context.
@@
Recent years have seen a flurry of advancements in generative artificial intelligence (AI) models, and their diverse applications [27; 28] have leveraged the potential of large-scale data via generative models, motivating the primary thesis of this work: generative models can learn the underlying mechanisms of traffic from low spatio-temporal resolution sensor input to generate enhanced data which better capturing stop-and-go waves, almost to the resolution of high-fidelity traffic sensors.
To investigate this, we collected 132 hours of rush-hour traffic data over 50 days from a stretch of 4 miles of roadway, amounting to over 2 million vehicle miles traveled (VMT). The dataset, named WaveX, was gathered simultaneously from two distinct sensor systems: one from a commonly used freeway radar detector system (with sensors deployed every 0.3 to 0.5 mile and aggregated every 30 seconds), and the other from a camera-based traffic testbed capable of capturing individual vehicle trajectories (aggregated into speed profile data with bin resolution of 4 seconds and 0.02 miles, approximately 32 meters) [17; 29]. We then formulated the data refinement problem as a traffic measurement super-resolution reconstruction task, with the goal of reconstructing the higher-fidelity data using only the lower fidelity data as input. To this end, we implemented a conditional diffusion denoising model for iterative refinement. Finally, we evaluated the extent to which our proposed method accurately reconstructs the details of stop-and-go waves.
@@
  1. (i)
    We are the first to apply a generative conditional diffusion denoising model designed to improve the spatio-temporal resolution of conventional freeway roadside traffic sensors, addressing the traffic measurement super-resolution challenge.
  2. (ii)
    We produce the WaveX dataset, a new large-scale empirical dataset over 2 million vehicle miles traveled featuring stop-and-go traffic, which was collected concurrently using both high-resolution and low-resolution sensor systems over 50 days, 4 hours a day. This dataset is intended to investigate the potential of generative models in analyzing traffic flow.
  3. (iii)
@@
  4. (iv)
    We release the dataset along with the code for community-wide use to enable future exploration of generative AI in transportation and support the reproducible research.
The remainder of this article is structured as follows: Section 2 provides a brief overview of related work on traffic waves and the use of generative models in traffic. Section 3 details the proposed method. Section 4 presents the collected data, while the experiments conducted to validate our approach are described in Section 5. The results are outlined in Section 6, and further discussed in detail in Section 7. Section 8 highlights the ability of our method to capture fine-grained wave properties and discusses its limitations.
@@ 2 Related work and challenges
Traffic reconstruction across all scales [30] requires data collection as the first step. Efforts in empirical traffic data collection have been ongoing since 1933 [31], and these efforts can be categorized into microscopic and macroscopic scales. Macroscopic traffic data [31; 32; 33; 10; 34; 8], which refers to the aggregated traffic measurements typically collected from loop detectors, provides insights into collective phenomena such as the fundamental diagram and capacity drop [35; 36]. On the other hand, microscopic traffic data such as vehicle trajectory data in a common reference system[37; 22], helps researchers understand individual vehicle behaviors.
@@
Traffic data which captures observations of both microscopic and macroscopic phenomena has significant value in the transportation research community. A relative scarcity of these data has limited empirical validation of theory in individual dynamics and traffic system dynamics. To date, only a few datasets are available which capture both microscopic and macroscopic data. Data collected from mobile and fixed sensors by the Mobile Century experiments [38] have been utilized for traffic reconstruction. Additionally, INRIX and loop detector data [13] have been collected and employed for cross-validation purposes. Purveyors of GPS-navigation applications have vehicle trajectory [39] and aggregate data, but these are proprietary. Further efforts and contributions are needed to expand the availability and diversity of datasets in this field.
@@
Ideally, microscopic and macroscopic scales in traffic measurement could be converted into another, similar to the concept of downsampling and upsampling in image processing [40]. In practice, transitioning from microscopic to macroscopic is straightforward through aggregation via Edie’s definition [2], whereas transforming from macroscopic to microscopic is more challenging due to the complex and heterogeneous dynamics of individual particles in the traffic flow system. The problem is often referred to as traffic reconstruction [38; 41; 42; 43]. One paradigm from previous research involves calibrating sets of microscopic modeled driving behaviors to replicate macroscopic measurements or phenomena [44]. A limitation of using micro-simulation to reproduce macroscopic phenomena is the inherent mismatch between the model and real-world aggregated data, which may not fully capture the complex and dynamic nature of actual traffic systems [45; 41]. Another paradigm involves using kernel smoothing [46; 47] and similar techniques to interpolate and impute sparse observations, in order to enhance the resolution and completeness of traffic data by filling in gaps and smoothing out inconsistencies. The adaptive smoothing method [48; 46; 49] is a widely recognized approach in this direction. More recently, the refinement problem was proposed and defined by [50], which utilized a simple linear regression approach that accounts for different traffic states to enhance traffic measurement resolution. The method proposed in that study achieves a 2 by 2 lift in resolution for both space and time each time, demonstrating the potential to improve resolution from 200 meters by 60 seconds to 50 meters by 30 seconds. In other fluid dynamical systems, deep learning has been explored and applied to the reconstruction problem. These approaches have shown promise in accurately capturing and reconstructing complex fluid dynamics [51; 52; 53], suggesting potential applicability to traffic flow systems [54] as well.
@@
In recent years, the field of image processing has seen significant advancements due to generative artificial intelligence. Generative models have enabled the transformation of low-resolution images into high-resolution ones via a process known as super-resolution [55]. This aligns well with the problem of enhancing traffic measurement resolution. The concept of generative deep learning gained prominence with the introduction of generative adversarial networks (GANs) [56], which have been extensively applied in traffic systems for tasks such as traffic data imputation [57; 58] and traffic state estimation [59]. More recently, diffusion probabilistic models have emerged as a promising approach [28]. These models have been investigated for tasks such as noisy data recovery [60] and GPS trajectory data generation [61], demonstrating their robustness to hyperparameters and their potential for a wide range of applications. These advancements in generative diffusion models offer new opportunities for improving traffic measurement resolution and addressing the challenges associated with low spatio-temporal resolution data, which have not been investigated before.
@@ 3.1 Problem formulation
In this article, stop-and-go wave refinement refers to the problem of recovering fine-grained speed data (high-fidelity) from corresponding coarse-grained traffic measurement data (relatively low-fidelity), which can be viewed as a super-resolution task. Coarse-grained data typically collected from radar detectors or loop detectors, is considered low spatio-temporal resolution data and is denoted by $\mathbf{r}\in\mathcal{R}$. Fine-grained speed data, generated from vehicle trajectory data, is viewed as high-fidelity data and is represented by $\mathbf{m}\in\mathcal{M}$. For a given space and time range, forming tensors $\mathbf{r}$ and $\mathbf{m}$ herein, each observation of coarse-grained data is represented as $\mathbf{r}\in\mathbb{R}^{T_{r}\times S_{r}\times 3}$ and fine-grained data is represented as $\mathbf{m}\in\mathbb{R}^{T_{m}\times S_{m}\times 1}$, where $T_{r},~{}S_{r},~{}T_{m},~{}S_{m}$ are the number of time and space observations in the space-time range, respectively. Typically, $T_{r}<T_{m}$ and $S_{r}<S_{m}$, which induces a dimension lifting in the output. The coarse-grained data contains three traffic measurements: speed, volume, and occupancy, whereas the fine-grained data includes only one measurement: speed. The problem addressed in this context is to find a mapping function $f:\mathcal{R}\to\mathcal{M}$ that can restore the fine-scale details of stop-and-go waves in traffic from the coarse-grained data.
@@ 3.2 Conditional denoising diffusion model
We adapt the denoising diffusion probabilistic model (DDPM) [62] for conditional fine-grained speed field generation. The method contains a noising process (e.g. forward diffusion) and a denoising process (e.g. inverse diffusion) as shown in Figure 2. These processes are defined as follows.
@@
$q(\mathbf{m}_{1:T}\mid\mathbf{m}_{0})$ $=\prod_{t=1}^{T}q(\mathbf{m}_{t}\mid\mathbf{m}_{t-1}),$ (1)
$q(\mathbf{m}_{t}\mid\mathbf{m}_{t-1})$ $=\mathcal{N}(\mathbf{m}_{t}\mid\sqrt{\alpha_{t}}\mathbf{m}_{t-1},(1-\alpha_{t})\mathbf{I}).$ (2)
@@
where $\gamma_{t}=\prod_{i=1}^{t}\alpha_{i}$.
Learn the inverse diffusion (denoise): to reverse the forward diffusion process, a neural denoising model $f_{\theta}$ takes the coarse-grained data $\mathbf{r}$ and the noisy target data $\mathbf{\tilde{m}}$ as defined in (4):
$\mathbf{\tilde{m}}=\sqrt{\gamma}\mathbf{m}_{0}+\sqrt{1-\gamma}\epsilon,~{}~{}~{}\epsilon\in\mathcal{N}(0,\mathbf{I}).$ (4)
The model $f_{\theta}(\mathbf{r},~{}\mathbf{\tilde{m}},~{}\gamma)$ is trying to learn from the coarse data $\mathbf{r}$, the noisy target image $\mathbf{\tilde{m}}$, as well as the variance of the noise $\gamma$, to predict the noise vector $\epsilon$ in (4). Here, $\gamma$ is the scalar parameter drawn from the distribution $p(\gamma)$. The objective function for training $f_{\theta}$ is written as:
@@
During the training process, each step involves performing a gradient descent update to minimize the loss function and optimize the model parameters:
$\nabla_{\theta}\left\|f_{\theta}(\mathbf{r},\sqrt{\gamma}\mathbf{m}_{0}+\sqrt{1-\gamma}\epsilon,\gamma)-\epsilon\right\|_{2}^{2},$ (6)
until convergence. The detailed proof and justification of the choice of the objective function for $f_{\theta}$ from the perspective of a variational lower bound and denoising score-matching can be found in [28; 62].
@@
$\mathbf{W}_{ij}(\mathbf{m},v_{c})=\begin{cases}\omega&\text{if }\mathbf{m}_{ij}<v_{c}\\ 1-\omega&\text{if }\mathbf{m}_{ij}\geq v_{c}\end{cases},$ (7)
@@
where $z$ is defined as Gaussian noise for all the steps other than $t=1$ (i.e., $z\sim\mathcal{N}(0,\mathbf{I})\text{ if }t>1,\text{ else }z=0$). The derivation of (9) can be found in [62].
Neural denosing model choice: the UNet [64] architecture is a popular choice for diffusion denoising model regression [28; 62] because its encoder-decoder structure with skip connections effectively captures multi-scale features and preserves high-resolution details throughout the denoising process. Detailed task specific architectural details can be found in Section 5.